OpenStudy (anonymous):

What is the simplest factored form of 4t2 − 16?

OpenStudy (turingtest):

http://openstudy.com/study#/updates/4f10409ae4b04f0f8a91f985 why are there two of the same problem?

OpenStudy (anonymous):

4(t+2)(t-2)

OpenStudy (wasiqss):

see, the 16 will go other side for simplifying then it equals 4t^2=16 we divide both sides by 4 to get x^2=4 hence as u can see two values of x will satisfy this equation that r +2,-2, hence x=+2, x=-2

OpenStudy (anonymous):

4t2 − 16=4(t2-4)=4(t+2)(t-2)